Yvettes Snack Shop sells two brands of potato chips. She produces them by buying them from a wholesale supplier. Brand X costs Yvette $1 per bag, and Brand Y costs her $1.40. Assuming Yvette has $280 budgeted to spend on the purchase of potato chips from the wholesaler, complete the following table showing her production possibilities. Brand XBrand Y (Bags) (Bags) i0 280 210 140 70 10 200 Now plot Yvettes production possibilities frontier in the following graph
Answer #1

(a) Budget line: $280 = $1X + $1.4Y

280 = X + 1.4Y

Y = (280 - X) / 1.4

Therefore:

Brand X Brand Y
280 (280 - 280) / 1.4 = 0
210 (280 - 210) / 1.4 = 70 / 1.4 = 50
140 (280 - 140) / 1.4 = 140 / 1.4 = 100
70 (280- 70) / 1.4 = 210 / 1.4 = 150
0 (280 - 0) / 1.4 = 200

(b) PPF as follows.

200 175 T 150 125 100 n 75 50 T 25 0 70 140 210 280 Brand X (Bags)

(c) Production possibilities frontier is not bowed out because principle of increasing costs is not applicable in this case (Since PPF is a straight line, it means constant opportunity costs).